OpenStudy (anonymous):

Solve this system {4a-3b=-3 6a-b=6}

OpenStudy (aripotta):

hello :) there's a few ways to solve systems of equations. which way are you most comfortable with? substitution method or elimination method?

OpenStudy (anonymous):

right, there are 2 ways to solve this.

OpenStudy (anonymous):

Um whichever way makes more sense. Im not really sure.

OpenStudy (aripotta):

ok, well let's do the substitution method.

OpenStudy (aripotta):

i'm going to use the second equation to solve for b 6a - b = 6 -b = -6a + 6 b = 6a - 6

OpenStudy (anonymous):

gotcha

OpenStudy (aripotta):

now i'm going to substitute that for b in the first equation

OpenStudy (aripotta):

4a - 3b = -3 4a - 3(6a - 6) = -3 4a - 18a + 18 = -3 -14a + 18 = -3 -14a = -21 a = 1.5

OpenStudy (aripotta):

now that we know what a equals, i'm going to substitute that in either equation

OpenStudy (anonymous):

okay

OpenStudy (aripotta):

6a - b = 6 6(1.5) - b = 6 9 - b = 6 -b = -3 b = 3

OpenStudy (aripotta):

so our answer is a = 1.5, b = 3

OpenStudy (aripotta):

make sense?

OpenStudy (anonymous):

part b is)

OpenStudy (anonymous):

thank you

OpenStudy (aripotta):

c would not work

OpenStudy (anonymous):

ok

OpenStudy (aripotta):

you'd have multiply the second equation by -4

OpenStudy (anonymous):

So its not C or D then

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

opps caught the "would not use"so its C

OpenStudy (aripotta):

well it asks which method would \(\underline{not}\) work

OpenStudy (anonymous):

gotcha sorry

OpenStudy (aripotta):

exactly, so c is your answer

OpenStudy (anonymous):

Could you help me with the next review problem? Please? Thanks!

OpenStudy (aripotta):

yes, sorry. i was helping someone else

OpenStudy (anonymous):

thats okay!

OpenStudy (aripotta):

oh matrix... i haven't used that in forever :x lemme google it real quick for a refresher

OpenStudy (anonymous):

lol ok

OpenStudy (aripotta):

lol i think you might have to open a new question and get someone else to help with this. i'm sorry :/

OpenStudy (aripotta):

it's been like a yr and a half since i used matrices

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

okay, what about this one?

OpenStudy (aripotta):

so long as i don't have to use matrices, i can help lol

OpenStudy (anonymous):

ok lol

OpenStudy (aripotta):

do you wanna use the substitution or elimination method

OpenStudy (anonymous):

subsitution

OpenStudy (aripotta):

alright. what should we do first, then?

OpenStudy (anonymous):

use the second equation to solve the first one?

OpenStudy (anonymous):

6a-6=b?

OpenStudy (aripotta):

use the second equation to solve for b in terms of a

OpenStudy (anonymous):

oops we already did this one. I meant {3x + 4y = 13 2x + y = 7}

OpenStudy (aripotta):

oh lols. it felt familiar

OpenStudy (aripotta):

ok! we're still going to use the second equation to solve for y in terms of x

OpenStudy (anonymous):

y

OpenStudy (anonymous):

*yes

OpenStudy (aripotta):

2x + y = 7 subtract 2x from both sides y = -2x + 7

OpenStudy (chosenmatt):

yup

OpenStudy (aripotta):

and now that we've solved for y, we're gonna substitute that for y in the first equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Sooo 3x + 4(7-2x) = 13

OpenStudy (aripotta):

yes and now you can solve for x

OpenStudy (anonymous):

x = 8/5?

OpenStudy (aripotta):

3x + 4(7 - 2x) = 13 3x + 28 - 8x = 13 -5x + 28 = 13 -5x = -15 x = 3

OpenStudy (anonymous):

Oh whoops okay

OpenStudy (anonymous):

so now solve for y?

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

9 + 4y = 13 4y = 4 y =1?

OpenStudy (aripotta):

yes :)

OpenStudy (anonymous):

so y = 1, x = 3

OpenStudy (aripotta):

yea

OpenStudy (anonymous):

What about this one?

OpenStudy (anonymous):

?

OpenStudy (aripotta):

i'm thinking

OpenStudy (aripotta):

i'm not sure how to do it

OpenStudy (anonymous):

ok

OpenStudy (aripotta):

are there choices? if there are, you can just plug them in to each choice to see which one is true

OpenStudy (anonymous):

nope there are no choices

OpenStudy (anonymous):

what about this one?

OpenStudy (aripotta):

wait i got the answer to that last one: a = 7, b = -2, c = 0

OpenStudy (anonymous):

ok cool!

OpenStudy (aripotta):

y = -12 x = 4

OpenStudy (anonymous):

okay! what about this different problem?

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

9?

OpenStudy (anonymous):

How did you figure that out?

OpenStudy (aripotta):

well i'm not 100% that it's correct

OpenStudy (aripotta):

maybe it's -9

OpenStudy (aripotta):

yea i think -9

OpenStudy (aripotta):

3A + 5B = -12 -9A - 15B = 6 trying to solve with the elimination method, we'd have to multiply the first equation by 3, which would give us 9A + 15B = -36 -9A - 15B = 6 doing that would cancel both variables, so no solution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about this one?

OpenStudy (aripotta):

ok so the slope-intercept form is y = mx + b where m is the slope and b is the y-intercept. so first, we're going to find the slope by using the slope formula:\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (aripotta):

then look at the graph to see what the y-intercept is

OpenStudy (aripotta):

seeing a dotted line means it's going to be either < or >, not \[\le ~or~\ge \]

OpenStudy (aripotta):

so have you figured out the slope?

OpenStudy (anonymous):

um no i havnt. not sure how.

OpenStudy (aripotta):

well it gives you two points, (0,2) and (-1,0), and i gave you the slope formula: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

2-0 / 0 - 1 = 2/-1

OpenStudy (aripotta):

\[\frac{ 2-0 }{ 0-(-1) }=\frac{ 2 }{ 1 }=2\]

OpenStudy (aripotta):

and, looking at the graph, what is the y-intercept?

OpenStudy (anonymous):

um....not sure.

OpenStudy (aripotta):

at which point does the line cross the y-axis?

OpenStudy (aripotta):

ok so the y-intercept is (0,2). so so far we have y (< or >) 2x + 2